for the series 2+4+8+16+.......... find the sum of 5 terms
Answer by checkley77(12569)

SERIES=AR^(N-1)=2*2^(5-1)=2*2^4=2*16=32
2+4+8+16+32
SUM=(A-RL)/(1-R) WHERE A=THE FIRST TERM, R=COMMON RATIO & L=LAST TERM.
SUM=(2-2*32)/(1-2)=(2-64)/-1=-62/-1=62 ANSWER.